Edit View Insert Format Tools Table Window Help nsert Design Layout References Mailings Review » share Calibri (Body) ,1111 | A-Av A-A2 Styles a the degree of relationship without regard to the torm of the relationship b. the degree to which the relationship is consistently multi-directional c. the degree of linear relationship d. the degree of curvilinear relationship A scatter plot shows a set of data points that are clustered close to a line that slopes down to the right. Which of the following values would be closest to the correlation for these data? a. 0.80 b. 0.40 c. -0.40 d. -0.80 11. 12. The results from a research study indicate that adolescents who watch more violent content on televisi also tend to engage in more violent behavior than their peers. The correlation between amount of TV violent content and amount of violent behavior is an example of a. a positive correlation. b. a negative correlation. c. a correlation near zero. d. a correlation near one. 13. A researcher reports that there is no consistent relationship between grade point average and the number of hours spent studying for college students. The correlation between grade point average and the numb of hours studying is an example of a a positive correlation. b. a negative correlation. c, a correlation near zero. d. a correlation near bne 14. A researcher conducts an experiment comparing three treatment conditions with t-10 scores in each condition. If the researcher uses a repeated-measures design, what are the df values for the F- ratio a. 1,9 b. 2,9 c. 2,18 d. 1,18 15. A researcher uses a repeated-measures ANOVA to evaluate the results from a research study and reports an F-ratio with df 2.30. Hlow many treatment conditions were compared in the study? b.3 c.4 d.s İfferen c e ba Independent and Renea red measure desiens is that repeated measure,
A rwannend%ct refers to differences between scores that are cause conditions, The differences (or variability) produced by treatment e a the numerator of the F-ratio b the denominator of the F-ratio c. both the numerator and the denominator of the F-ratio d. Theatment effects do not contribute to the F-ratio because they an ratio is computed. he Guermine which reatinents ur lifeni b. determine how much difference there is between treatments c. determine whether or not a Type I Error was made inthe ANOVA d. determine whether or not a complete ANOVA is justified 6. In analysis of variance, MS provides a measure of_. a. variance b. average differences among means c. the total variability for the set of N scores d. the overall mean for the set of N scores 7. In analysis of variance, each independent variable is referred to as a a. factor b. level c. comparison d. stage A positive value for a correlation indicates 8. increases in X tend to be accompanied by increases in Y a. b. increases in X tend to be accompanied by decreases in Y a much stronger relationship than if the correlation were negative a much weaker relationship than if the correlation were negative c. d. 9. A negative value for a correlation indicates
media%2Fa7d%2Fa7dda602-453c-496a-9082-a9
Answer #1

For slide 1 question 14-

Given that there are three treatments and n =10

So for treatment degrees of freedom are=k-1

Where k = Total no. Of treatment

=3-1

=2

For error degrees of freedom= (k-1)*(n-1)=2*9=18

Therefore for df for f ratio is 2,18

For slide 2 questions 7

In analysis of variation. Each independent variable referred to as factor.